### 15 Dec 2023 v1

# ACPO: AI-Enabled Compiler-Driven Program Optimization

Abstract.
The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler. Traditionally, such profitability decisions are made by hand-coded algorithms tuned for a very small number of benchmarks, usually requiring a great deal of effort to be retuned when the benchmark suite changes. This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework. Even so, seamless integration of ML into the compiler would require constant rebuilding of the compiler when models are updated.
This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes. We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes. Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% and 3%, 5.4%, 0.2% compared to LLVM’s O3 optimization when deployed on Polybench, Coral-2, CoreMark, and Graph-500, respectively. Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
Compilers, Optimization Framework, Loop Unrolling, Inlining, Deep Learning, LLVM
copyright: nonejournal: tacoccs: Software and its engineering Compilersccs: Computing methodologies Machine learning{textblock}
13(1.6,1) Under revision at ACM Transactions on Architecture and Code Optimization (TACO)
Received 26 June 2023; Revised 20 Nov 2023
LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
1. Introduction
State-of-the-art compilers are built with sophisticated optimization pipelines that result in robust, reproducible (Ivie and Thain, 2018), and reliable code generation (Aho et al., 1986; Hall et al., 2009). The LLVM compiler framework is no exception with its multi-layer design, where passes in the mid-end can perform analyses or transformations on a code region, one after another (Lattner, 2008). LLVM passes operate on code regions with varying levels of granularity, including loops, functions, call graphs, and modules. Compiler engineers have continuously improved the quality of the optimization passes in LLVM, and have derived a set of fixed-length optimization pipelines, known as standard optimization levels, i.e., O{1,2,3,s,z}, which are designed to improve the code size or performance (to varying degree) of the programs to which they are applied. At the highest optimization level available, i.e. O3, more than 60 unique transformation passes exist in LLVM’s pipeline (Ashouri et al., 2017).
Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years. In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023). Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set. Alternatively, by the growth of applications of Reinforcement Learning (RL), as opposed to generating offline training data by an autotuner, authors have explored their action space, online, and during the training process of the RL agent. This is done by means of feeding back a reward value to update the agent’s policy (Trofin et al., 2021; Cummins et al., 2021b; Ashouri et al., 2022; Wang et al., 2022a). To characterize the software being compiled for the ML model, program representation methods are employed, which generally fall into three categories: using static features (Fursin et al., 2011; Ashouri et al., 2022), using dynamic features (Cavazos et al., 2006; Hoste and Eeckhout, 2007), and graph-based characterization (Park et al., 2012; Ben-Nun et al., 2018; Alon et al., 2019; Cummins et al., 2021a). In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
Recently, there have been a few works providing an end-to-end approach to construct ML-based heuristics for optimizing the decision making process of the transformation passes (Cummins et al., 2017; Das et al., 2020; Rotem and Cummins, 2021; Trofin et al., 2021; Cummins et al., 2021b; Mannarswamy and Das, 2022; Ashouri et al., 2022). Among these, only a subset provides an ML-guided optimization approach (Cummins et al., 2021b; Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022) baked into a single or multiple passes to readily streamline the optimization process inside the pass functionality. The recent advancements in this area beg the question of how scalable, user-friendly, and comprehensive a framework should be, to provide seamless integration with existing LLVM optimization passes, while maintaining a clean separation of concerns between the ML and the LLVM layers.
The vision for ACPO is to enable the use of Machine Learning in a compiler, in a way that is modular, extendable, and easy to use. ACPO aims to allow for easy collaboration between compiler developers and ML architects by use of clear and simple APIs that enable the collection of relevant data for training purposes, providing inputs to an ML model during inference, and separating the compiler from the model and ML framework to enable model development and testing within the compiler context with minimal intrusion into the compiler code base. This enables the compiler engineers to optimize passes with an easy-to-use set of tools and limited ML knowledge. At the same time, once an ML model is integrated into the compiler, it can be easily optimized and redesigned with no changes to the compiler code base, effectively enabling ML architects to contribute to compiler improvement. This is possible due to ACPO’s separation of the ML framework and the compiler, making it possible for the compiler to work just as easily with any ML model as it is to work with any ML framework, e.g., TensorFlow (Abadi et al., 2016), PyTorch (Paszke et al., 2017), etc, and even other representation mechanism (Fursin et al., 2008; Hoste and Eeckhout, 2007; Cummins et al., 2021a). To this end, ACPO proposes the following contributions:
  1. (1)
    As a framework, it provides a comprehensive set of tools, libraries, and algorithmic methods by enabling compiler engineers with a straightforward and user-friendly interface to instantiate ACPOModel classes to replace LLVM’s existing hard-coded heuristics.
  2. (2)
    ML APIs and the compiler are seamlessly connected, but at the same time, they are not interdependent; changing compiler versions, ML models, or ML frameworks won’t break the functionality of ACPO. As long as the inputs/outputs match, users can easily revise the ML side without the need to rebuild LLVM after every change.
  3. (3)
    Showcasing the benefits of ACPO, we demonstrate two different scenarios:
    1. (a)
      Loop Unroll Pass — Building both the interface and the ML model. Loop Unroll pass can be optimized from scratch by adding the necessary code to instantiate ACPOLUModel from ACPOModel and also by providing an ML model to outperform LLVM’s unrolling decisions at O3 on Polybench and a number of other benchmarks.
    2. (b)
      Function Inlining Pass — Building the interface and leveraging an existing ML model. Function Inlining pass can be optimized by leveraging an existing ML model, i.e., MLGOPerf (Ashouri et al., 2022), using ACPO framework. We do that by creating an ACPOModel instance, called ACPOFIModel, and ensuring its inputs and outputs match the existing model so it can be simply invoked at the appropriate point in the compilation process. We demonstrate the ease of integration and versatility of ACPO framework and that how it can be used to connect existing models to the new frameworks. Experimental results on Cbench reveal that the performance of ACPOFIModel is comparable with the reported results in the original work while benefiting from a multi-layer design of ACPO on the separation between the ML APIs and the compiler.
The rest of the paper is organized as follows. Section 2 discusses the state of the art. Section 3 gives details on our proposed framework and how the different layers of abstraction communicate with each other. Under Section 4.1, we showcase how Loop Unroll pass can benefit by instantiating ACPOLUModel and, subsequently, Section 4.2 does so for Function Inlining pass. Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
2. Related Work
A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018a) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics. These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a). Further classification of the two aforementioned problems and by the type of the approach, provides insights for categorizing the works into three groups: (1) autotuning-based approaches (Fursin et al., 2011; Chen et al., 2012; Ashouri et al., 2013; Ansel et al., 2014; Kalyan et al., 2019; Cummins et al., 2021b; Wang et al., 2022b; Park et al., 2022), (2) end-to-end frameworks of optimizations (Fursin et al., 2008; Cummins et al., 2017; Haj-Ali et al., 2020; Cummins et al., 2021b; Rotem and Cummins, 2021; Das et al., 2020; Mannarswamy and Das, 2022; Ashouri et al., 2022; Wang et al., 2022b), and (3) ML-guided approaches (Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022).
Autotuning-based approaches leverage an iterative methodology to wrap around the compiler and to speed up the search in finding better solutions to optimizing a single or multiple passes, often constructing a new sequence of optimization to outperform standard optimization levels by means of phase-ordering or optimization selection (Agakov et al., 2006; Ashouri et al., 2016; Cummins et al., 2021b; Wang et al., 2022a). This can be done in a coarse (Kulkarni and Cavazos, 2012; Ashouri et al., 2016, 2017) or fine-grained manner (Kalyan et al., 2019; Phothilimthana et al., 2021; Cummins et al., 2021b). End-to-end approaches often provide a framework that can be paired or built with the compiler and to offer boosted search or ordering of optimizations (Cummins et al., 2017; Haj-Ali et al., 2020; Mannarswamy and Das, 2022). ML-guided approaches build the decision-making capabilities of ML models into single or multiple passes that can act as a standalone ML-enabled pass. Their corresponding ML models can be either built with the compiler using an Ahead-of-Time (AOT) compilation or be used at runtime using an alternative interprocess communication method, e.g., Python named pipes. All things considered, several recent works can be included in multiple classes, therefore, the lists here are not mutually exclusive.
Cummins et al. (Cummins et al., 2021b) propose CompilerGym, a fine-grained RL environment to speed up the search for tuning a single or multiple compiler passes defined as action space. The authors leverage OpenAI’s Gym (Brockman et al., 2016) library to provide a flexible environment to tackle the selection and the phase-ordering of optimization. The results showcase the benefits of using their RL agent to further reduce the code size of Cbench benchmarks. Contrary to this work, we are proposing an ML-guided approach to provide a framework to optimize passes. Additionally, our objective is to improve the performance as opposed to reducing the code size of a binary.
Rotem and Cummins (Rotem and Cummins, 2021) propose a Decision Trees (DS)-based framework to infer branch probabilities for use in providing profile-guided optimizations (PGO). The authors use hand-crafted features for the characterization of code regions and show they are able to reap the benefits of PGO without actually running the program with instrumentation to collect profile data and recompiling. This work is relevant to ours since the approach builds into LLVM and once trained, it can provide PGO certain profiling information without actually running PGO.
Trofin et al. (Trofin et al., 2021) propose MLGO that uses ML to advise the Inliner as to whether or not a call site should be inlined, in order to minimize the size of the generated code. MLGO framework compiles the ML model into C++ libraries using Ahead-of-Time (AOT) compilation which becomes part of the compiler at build time. Additionally, the authors have extended the code base to provide optimization for the register allocation problem. Our previous work, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g. SPEC CPU 2006 and Cbench.
In this work, we focus on the hard problem of ML performance optimization, where the key challenge is the development of model architecture and the features used to effectively capture key performance use cases and to facilitate the generation of robust ML models. In the process, we develop a scalable, extendible, end-to-end ML-guided framework to enable easy use of ML within optimization passes in compilers such as LLVM. We demonstrate the effectiveness and the generality of our framework by applying ML-enabling Loop Unroll and Function Inlining passes. We show how ML can easily and effectively be incorporated into a compiler while maintaining separation from the underlying ML frameworks. This framework is easily applicable to other passes and tasks within the compiler domain.
3. Proposed Methodology
Figure 1 presents the high-level design of our infrastructure. There are multiple components involved with the ACPO architecture: Autotuner, compiler, ACPOModel abstraction, low/high-level ML APIs, and ML Framework. For the most part, the autotuner is a wrapper around the compiler and its hooks in the compiler are used to generate model training data. During the inference flow, i.e., when ML models are used for decision making within the compiler, the autotuner is not used and we can think of the compiler as being at the top of the stack.
Figure 1. ACPO Architecture Infrastructure
The ACPO inference flow begins when an ML-enabled pass is invoked and creates an instance of ACPOModel. The instance of ACPOModel aggregates appropriate features from the code region being analyzed and specifies the kind of output required from the ML model. The ACPOModel object then transfers the features, along with information about the model being invoked, to our ML framework via a set of remote procedure call APIs. The APIs provide the ability to send requests, including loading an appropriate model into the ML framework, executing an inference call, and returning the result of the inference back to the compiler. The compiler then applies transformations to the input program as prescribed in the inference output it receives.
In the following sections, we describe in more detail the training and inference flows, and then elaborate on each component of the ACPO framework to show how they fit with the training and inference flows.
(a) Training Flow
(b) Inference Flow
Figure 2. ACPO Flow
3.1. Autotuner
An autotuner is a software tool that derives decisions for an optimization scenario (minimization or maximization), by intelligently exploring the design space of compiler optimization parameters (Palermo et al., 2005; Ansel et al., 2014; Ashouri et al., 2018a, b). The parameters can be either coarse- or fine-grained. The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g. instructions, loops, functions, etc., independently of each other. In this work, we perform fine-grained autotuning with an autotuner (Kalyan et al., 2019), which is built on top of OpenTuner (Ansel et al., 2014). This allows us to generate meaningful training data and target labels that capture the effect of optimization parameters on code region types of interest.
3.2. ACPO Training
A key aspect of creating an ML-driven compiler is the training of models on sample data to ensure that the compiler makes decisions consistently with good results contained in the training data. To do this, we opted for a two-step approach: generate data based on a set of benchmarks the model is expected to reasonably reproduce and then train the model with the collected data. The data generation step is shown in Figure 2a.
To generate data for training purposes we rely on the autotuner. The autotuning process allows us to override compiler heuristics with an iterative search for the optimal optimization decisions for a given application. By performing the search on our benchmark programs and recording each step of the search, we can obtain information about the impact of each decision on program performance, and thus collect sufficient information to be able to train the eventual model with a reasonable chance of success. To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another. For example, in the case of loop unrolling, we enabled the autotuner to search for a good unroll factor for each loop, instead of applying an unroll threshold that would be specified on the compiler command line and hence applied globally on all loops. With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks.
The first step in the autotuning flow is to establish a search space over which the tuning occurs. This comprises both the focus on a pass as well as code regions on which the specific transformation can act. In loop unrolling for instance, the applicable code regions would be loops, not single instructions or whole functions. Once the search space is established, the autotuner simply proposes solutions for each code region and then requests the compiler to apply the desired optimization on the given code region at the appropriate time. This is crucial because we do not want to affect the compiler pipeline itself; instead, we want to generate a wide array of solutions at a specific point in the compilation process and then observe the impact of those decisions. That is key to creating a viable model.
At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions. Initially, the pass proceeds normally, starting with legality checks and then following through to profitability analysis, which is when the ML model will be invoked. The compiler instantiates an ACPOModel and gathers features pertaining to the code region being optimized. This is important to ensure that the model being trained is provided with input features based on what they will be at deployment time when the compiler uses this ML-enabled pass. Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation. The features are passed with std::vector of pairs to the ML side. This process can also be scaled to an arbitrary number of features and feature types. The feature set is then saved for later use to allow the features, the decision specified by the autotuner, and the final impact of the decision to be correlated. The transformation decision as prescribed by the autotuner is applied to the code region (e.g. the autotuner can set the unroll factor of a loop to a specific value), and then the compilation continues normally until it reaches another code region that requires a decision from the ML model, at which time the process is repeated.
Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program). The measurement is used as an assessment of decision impact, so the decisions that lead to a positive impact can be distinguished from the ones that lead to a negative impact on performance. Generally, we use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation. This process is repeated many times to gather as much data per benchmark as possible to achieve beneficial results on most benchmarks. The resulting trial data is then processed and used as input to the model training.
3.3. ACPO Inference
The compilation process involving ML inference is depicted in Figure 2b. The input program is supplied to Clang to be translated into LLVM Intermediate Representation (IR). The LLVM IR is then passed through a sequence of passes and eventually reaches a pass that uses ACPOModel for its profitability analysis. The pass proceeds normally to confirm the legality of the transformation; when it needs to decide the profitability of the transformation, the ML framework is engaged.
The first step in the inference process is the instantiation of the ML model, via the ACPOModel abstraction. This step sets up the interface to the underlying ML framework, such as TensorFlow (Abadi et al., 2016). Upon successful framework initialization, ACPOModel collects features required for inference and sends them to the ML framework. The ML framework is then instructed to run inference on the model, and the inference result is returned to the compiler. If the inference result is such that the transformation should be applied to a given code region, the transform pass then continues and applies the transformation. Otherwise, the pass gracefully rejects the option to transform the code and proceeds to either the next code region to be analyzed or terminates. In some cases, the inference result can contain a numeric value, e.g. the predicted optimal unroll factor for a given loop. To accomplish the tasks of training and inference we used several key components, which we describe in greater detail in the following subsections.
3.4. Compiler
The compiler of choice is LLVM (Lattner, 2008) version 12, which at a high level comprises the C/C++ frontend Clang, the mid-end optimizer, and a target-specific backend to generate executable code. The compiler first translates a program into LLVM IR, and then proceeds to optimize the IR via a sequence of passes. The passes are applied in sequence and at various scopes (module, function, loop, basic block). Once a pass has finished transforming a code region, appropriate internal data structures, including IR, are updated and the pass either proceeds to the next code region or terminates to enable the next pass to proceed. Once the compilation is done, an executable is produced that can be run and measured. We generally refer to the sequence of passes as a pipeline, and we focus on keeping the pipeline unchanged. This is important because empirical evidence shows that the order of passes and composition of the compiler pipeline affect program performance (Ashouri et al., 2017). In this work we focus on improving decision making within a pass using ML, so changes to the pipeline ordering and optimization selection of passes are orthogonal and outside the scope of the work.
One of the key things we added to the compiler is the ability to invoke a model and run inference given a set of features. To facilitate this, ML-enabled passes in the compiler have been given access to ACPOModel class, from which a pass-specific model instance can be derived.
3.5. ACPOModel
The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference. The header file snippets of ACPOModel can be found in the Appendix section under Listing 7.
This class is a simple abstraction of how the model is incorporated into the compiler flow. The class is instantiated and a reference to the ML Interface (ACPOMLInterface class) is provided to enable the class to communicate with the ML framework. This framework can be active in the background (as is the case in our implementation) or be created on an as-needed basis. In either case, the reference to the interface is provided to establish communication. The model is then configured to return a specific type of result as a part of the ACPOAdvice class. Then the features to be used as input for inference are specified as input to addFeature method. Finally, to perform inference the caller would invoke the getAdvice() method and act on the result provided.
The getAdvice() method takes several simple steps, most importantly, does prepare the input data by first invoking prepareModelInput() to ensure that model inputs are provided to the model correctly and then invoking the runModel() method. The runModel method establishes communication with the ML framework, loads the model if it is not yet loaded, sends input features in, and then requests that the inference be run. If at any point in time there is an error, the process is aborted and the runModel() method will return false, subsequently causing ACPOAdvice to set to nullptr. This will be a crucial check to verify that ML inference is completed and should be checked by the caller.
It is worth noting that the prepareModelInput() and runModel() methods are virtual. The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
3.6. ML APIs
An important part of the ACPOModel class is the reference to the ACPOMLInterface class. This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler. The snippet for the ACPOMLInterface header is shown under the Appendix section in Listing 8.
The ACPOMLInterface class is virtual and it is meant to be implemented depending on how your particular ML framework needs are accessed. If it is a compiled library linked with your version of LLVM or a separate process (as was the case for this paper), this class abstracts the exchange of data between the compiler and the ML framework so that it is fundamentally irrelevant to how the ML framework actually works. To use this class, the ACPOModel invokes routines such as loadModel, freeModel, setCustomFeatures, and others to set up the model on the side of the ML framework. Then when the input data is ready, runModel is invoked to perform inference. The result is then retrieved by the ACPOModel via the getModelResult family of methods.
In our implementation, we created a derived class that communicates with an ML framework using named pipes, i.e. os.mkfifo(). The ML framework in our scenario is a separate python process that has a user-visible interface through the named pipes. It receives commands from ACPOMLInterface-derived class on the compiler side, processes the requests, and responds with requested information or status. This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis. While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes. We envision this will grow to comprise a toolkit with many different models, algorithms, and program representation mechanisms at the developer’s disposal. Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 3.
3.7. ML Framework
In this work, to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment. Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects.
4. ACPO Models
In this section, we provide two use cases of leveraging ACPO framework on optimizing compiler passes. As we mentioned earlier, state-of-the-art approaches showcase the benefits of adapting ML models for coarse-grain and fine-grain tuning of a number of compiler optimization passes We speculate, ACPO can benefit any mid/back-end optimization/transformation pass with an internal decision process mechanism and a set of inputs/outputs. Historically, compiler developers have incrementally improved such optimization passes by means of algorithmic and heuristic approaches (Aho et al., 1986; Hall et al., 2009) and we believe leveraging ML to enhance the internal decision-making process of optimization passes merits an in-depth investigation and lays the groundwork for many future works. To this end, we choose Loop Unroll and Function Inlining as two of the most important compiler optimizations used in many state-of-the-art compilers, including LLVM.
4.1. ACPO Loop Unrolling Model
Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018a). There have been multiple works trying to automatically construct heuristics to find the optimal unrolling factor (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018) and research has shown identifying the optimal unroll count is a fine line between the increase in speed versus the binary code size by which the increase in cache misses can be counterproductive. Therefore, a number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.
1class ACPOLUModel : public ACPOModel {
2public:
3  ACPOLUModel(Loop *L, OptimizationRemarkEmitter *ORE, bool UseML = true)
4      : ACPOModel(ORE, UseML), CurrentLoop(L), UnrollingFactor(0), UnrollingType(0) {
5        std::vector<BasicBlock *> BBVec = L->getBlocksVector();
6    ...
7    setMLIF(createPersistentPythonMLIF());
8  }
9  ...
10protected:
11  // Get an Advice inferred from the ML Model
12  virtual std::unique_ptr<ACPOAdvice> getAdviceML() override {
13    std::shared_ptr<ACPOMLInterface> MLIF = getMLIF();
14    MLIF->loadModel("model-lu.acpo");
15    MLIF->setCustomFeatures("LU", CustomFeatureValues);
16    bool ModelRunOK = MLIF->runModel("LU");
17    ...
18    UnrollingType = MLIF->getModelResultI("LU-Type");
19    UnrollingFactor = MLIF->getModelResultI("LU-Count");
20    ...
21    // Form Advice object to return to caller
22    std::unique_ptr<ACPOAdvice> Advice = std::make_unique<ACPOAdvice>();
23    Advice->addField("LU-Count", ConstantInt::get(Type::getInt64Ty(*(getContextPtr())),(int64_t)UnrollingFactor));
24    Advice->addField("LU-Type", ConstantInt::get(Type::getInt64Ty(*(getContextPtr())),(int64_t)UnrollingType));
25    return Advice;
26  }
27  ...
28};
Listing 1: ACPOLUModel Instantiation
A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1. We start by providing the necessary arguments for the constructor and initializing the private members. Subsequently, we run a few checks to make sure the steps mentioned prior under Section 3.6 are performed correctly. That leads us to the main inference method which is getAdviceML(). This method first retrieves a reference to the ML interface which is used to communicate with the ML framework. It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file. If the model has already been loaded by the framework, then this command completes instantly, as loading the same model many times serves no purpose. The method then sets the inputs to the model and invokes the runModel method to begin inference.
The inference process takes place on the ML framework, and the runModel call is a blocking call that waits until the inference completes. When it does, a successful inference will return desired results, which can be retrieved using getModelResultI methods. However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed. This mechanism allows the compiler to handle this case as the developer chooses. In our case, we forced the compiler to terminate with an error to ensure our ML model was the only mechanism for determining profitability. However, developers are welcome to provide fall-back mechanisms as they see fit.
To leverage ML in unrolling, we invoked the ACPOLUModel after the loop unrolling legality checks had been made. Also, we respected the high-priority overrides. The highest level is set by the explicit use of unroll-count by the user, followed by the usage of #pragma clang loop unroll_count(N). We invoke ACPOLUModel object subsequent to these two priorities and provide the unrolling information to the pass as shown in Listing 2.
1  if (EnableACPOLU.getNumOccurrences() > 0 && LAA) {
2    std::unique_ptr<ACPOLUModel> LU = std::make_unique<ACPOLUModel>(L, ORE);
3    ...
4    LU->setCustomFeatures(...);
5    ...
6    std::unique_ptr<ACPOAdvice> Advice = LU->getAdvice();
7    Constant *ValType = Advice->getField("LU-Type");
8    Constant *Val = Advice->getField("LU-Count");
9    ConstantInt *IntVal = dyn_cast<ConstantInt>(Val);
10    ConstantInt *LUType = dyn_cast<ConstantInt>(ValType);
11    ...
12  }
Listing 2: ACPOLUModel Invocation
In the code above we instantiate an ACPOLUModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework. The persistent interface is simply a parallel process that is launched when inference is required. However, unlike invoking blocking calls, the ML interface remains live as a separate process with which the compiler can communicate. In the case of a subsequent loop unrolling event, the process is not created anew. Instead, a reference to an existing interface is returned and the loop unrolling pass communicates with the same ML framework instance as in the previous inference step. This permits us to keep models live in memory through the compilation process. This is particularly useful in cases such as unrolling, which is often used more than once in a compiler pipeline. This way we do not have to pay the compile-time cost of restarting and closing the ML framework, or loading ML models into memory many times.
Once the ACPOLUModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOLUModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into unroll-type and unroll-factor fields, which are then obeyed by the unrolling pass and applied as directed. Note that in this work we only replace the profitability checks with ML models. Decisions regarding transformation legality are left with LLVM’s legality checks in a given pass, and if those checks fail, the transformation will not proceed, nor will the profitability model be invoked.
4.2. ACPO Function Inlining
In this section, we show how to leverage ACPO to instantiate ACPOModel class for function inlining. Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers. In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018a; Theodoridis et al., 2022). A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
1class ACPOFIModel : public ACPOModel {
2public:
3  ACPOLFIModel(CallBase *CB, InlineAdvisor *IA, OptimizationRemarkEmitter *ORE,
4               bool OnlyMandatory, bool UseML = true)
5      : ACPOModel(ORE, UseML), CurrentCB(CB), NotACPOAdvisor(IA), OnlyMandatory(OnlyMandatory) {
6    Function *Caller = CB->getCaller();    ...
7    setMLIF(createPersistentPythonMLIF());
8  }
9  ...
10protected:
11  // Get an Advice inferred from the ML Model
12  virtual std::unique_ptr<ACPOAdvice> getAdviceML() override {
13    std::shared_ptr<ACPOMLInterface> MLIF = getMLIF();
14    MLIF->loadModel("model-fi.acpo");
15    MLIF->setCustomFeatures("FI", CustomFeatureValues);
16    bool ModelRunOK = MLIF->runModel("FI");
17    ...
18    ShouldInline = MLIF->getModelResultI("FI-ShouldInline");
19    ...
20    // Form Advice object to return to caller
21    Advice->addField("FI-ShouldInline",
22                     ConstantInt::get(Type::getInt64Ty(*(getContextPtr())),
23                                      (int64_t)ShouldInline));
24    return Advice;
25  }
26  ...
27};
Listing 3: ACPOFIModel Instantiation
To leverage ML in function inlining, we invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made. Also, we respected the high-priority overrides. We invoke ACPOFIModel object and provide the inlining information to the pass as shown in Listing 4. We instantiate an ACPOFIModel instance, which in itself creates a persistent ML interface instance to a Python-based ML Framework, similar to what we showed earlier with loop unroll pass. Once the ACPOFIModel is instantiated, we invoke setCustomFeatures() to specify features as inputs to the model. This fulfills the requirements of the ACPOFIModel class and ensures all inputs are provided as needed before an inference is called. The getAdvice method is then called, which in turn invokes getAdviceML which facilitates the process of inference on behalf of the compiler. The result is then decomposed into ”FI-ShouldInline” field, which is then obeyed by the function inlining pass and applied as directed. Note that in this work, similar to loop unroll pass, we only replace the profitability checks with ML models.
1 /// helper function for getting advice with acpo infrastructure
2 bool getACPOAdvice(CallBase *CB, std::unique_ptr<ACPOFIModel> &FI,
3                   ModelDataFICollector *MDC, InlineAdvisor *Advisor) {
4  bool ShouldInline = false;
5  if (EnableACPOFI.getNumOccurrences() > 0) {
6    std::unique_ptr<ACPOFIModel> FI = std::make_unique<ACPOFIModel>(*CB, ORE);
7    ...
8    FI->setCustomFeatures(...);
9    ...
10    std::unique_ptr<ACPOAdvice> Advice = FI->getAdvice();
11    Constant *ValType = Advice->getField("FI-ShouldInline");
12    ConstantInt *ACPOInline = dyn_cast<ConstantInt>(Val);
13    ShouldInline = ACPOInline->getSExtValue();
14    return ShouldInline;
15  }
Listing 4: ACPOFIModel Invocation
5. Experimental Evaluation
In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of (1) standalone prediction accuracy, (2) cross-validation method using Polybench, and (3) the generalization evaluating of the model using a number of other standard benchmarks. Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
5.1. ACPO Loop Unrolling Model
As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side. Additionally, a new ML model has been designed and deployed for this matter. The details of the design and testing are provided as follows.
5.1.1. ACPOLUModel Architecture
Table 1 presents the architecture of ACPOLUModel. It is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor. i.e., UP.Count values, given a loop. We have designed and implemented our model in PyTorch (Paszke et al., 2017), but once we train it, we convert the model into a TensorFlow (TF) (Abadi et al., 2016) frozen model and leverage TF’s inference mechanism to do inference at deployment. As mentioned earlier in Section 3.6, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match.
Table 1. ACPO Loop Unrolling Model Architecture
Layer No
Layer (type) Output Shape
0
Linear-FC1 [Features, 1, 32]
Relu-0 [Features, 1, 32]
1
Linear-FC1 [-1, 1, 128]
Relu-1 [-1, 1, 128]
2
Linear-FC2 [-1, 1, 256]
Relu-2 [-1, 1, 256]
3
Linear-FC3 [-1, 1, 64]
Relu-3 [-1, 1, 64]
4
Linear-FC4 [-1, 1, Classes]
Softmax [-1, 1, Classes]
The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed. The features are passed with std::vector of pairs to the ML side. Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR. The features are shown in Table 2. The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e. classes corresponding to unrolling count, is depicted in Figure 3. The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label. We did this by using features with at least a $\pm$5correlation, and this can be seen in the final column/row. We collect these features prior to the actual loop unrolling transformation routine and leverage setCustomFeatures() method to communicate with the ML side via the MLInterface API. Note that feature engineering and software characterization (Fursin et al., 2011; Ding et al., 2019; Cummins et al., 2021a) is a dynamic and fast-paced topic in the space and is outside the scope of this work. We leave the usage of other methods for our future work.
Table 2. ACPOLUModel Features
No Static Features No Static Features
1
PartialOptSizeThreshold
16
NumLoadInstPerLoopNest
2
AllowRemainder
17
NumStoreInstPerLoopNest
3
UnrollRemainder
18
TotLoopNestInstCount
4
AllowExpensiveTripCount
19
AvgNumLoadInstPerLoopNest
5
Force
20
AvgNumStoreInstPerLoopNest
6
TripCount
21
NumLoadInstPerLoop
7
MaxTripCount
22
NumStoreInstPerLoop
8
Size
23
TotLoopInstCount
9
InitialIVValueInt
24
AvgNumLoadInstPerLoop
10
FinalIVValueInt
25
AvgNumStoreInstPerLoop
11
StepValueInt
26
IsInnerMostLoop
12
NumPartitions
27
IsOuterMostLoop
13
IndVarSetSize
28
MaxLoopHeight
14
AvgStoreSetSize
29
TotBlocksPerLoop
15
AvgNumInsts
30
IsFixedTripCount
Figure 3. Features Correlation Analysis
5.1.2. Data Collection
We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are $\in\{0,2,\cdots,64\}$. The autotuner uses the list of loop unrolling count opportunities to determine optimal configurations by using different search techniques. Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file. We run the emitted executable and provide the execution times to the autotuner as feedback which will be leveraged by the autotuner to generate a new set of configurations based on the feedback. This process was repeated for 100 iterations to generate our training data. Note that each iteration, depending on the number of loops an application has, generates multiple data points, hence the total number of training data is significantly higher than the number of iterations of tuning rounds. The data generation process takes around 2 days.
The execution time of a number of Polybench applications can be too short to allow for a reliable measurement and to this end, we define a larger input dataset for those having an execution time of less than a second. This includes: gemver, jacobi-1d, bicg, atax, mvt, durbin, trisolv, gesummv, and deriche. In this work, we define the Loop Unrolling speedup as the following:
(1) $\footnotesize Loop_{speedup}=\frac{Loop_{runtime(Base)}}{Loop_{runtime(X)}}$
where $Loop_{runtime(Base)}$ is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner. In this work, total program runtime and the $T_{Loop}$ are collected with Linux perf. All the runtime measurements are done single-threaded, and we use numactl tool to bind the workloads to a unique CPU core. All experiments are measured on an ARMv8.2-A (Kunpeng 920) architecture @ 2.6GHz on Linux. We run each benchmark three times and before each measurement run, we flush the system page cache to avoid any perturbance into the collection of our training data and we repeat the process for all measurements having more than 1% of variance between the runs to make sure they are robust.
Additionally, we treat a single instance of iteration of data generation, normally having multiple loops and generating multiple samples, one set of runs meaning that we assign the Loop_speedup value to the whole set of rows generated as a result of the configuration. This way, we can filter out bad runs from good runs simply by pruning those causing a slowdown, or even having at least a certain amount of speedup values. This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
5.1.3. Analysis of Compilation Time Overhead
We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table 3. As mentioned earlier under Section 4.1, Loop Unroll pass is registered twice at O3’s pipeline. We have implemented ACPO in a way that we load our ML models once per each module and keep the models alive in the lifetime of a module to avoid a repeated loading process of the models per each instance of inference. To this end, the other contributing factor is the number of modules per application and to a lesser extent the number of loops inside a module. The number of modules determines how many times the ML model needs to be loaded, which is an expensive task, and the number of loops determines the number of times Loop Unroll pass will be activated and ACPOLUModel will be deployed. ACPOLUModel is persistent and loaded once per module and we are working on making it even more optimized, i.e., loaded once per each application having multiple modules. For example. On Polybench’s bicg, we have two modules, namely, polybench.c and bicg.c. The former has a single loop and the latter, has 8, thus, the overall overhead would be 0.857 seconds.
Table 3. Compile-time Overhead Breakdown
Task Each Module Each Loop
Feature Collection 0.000135 0.000135
Setting Features for ML Model 0.000029 0.000029
Total ML Inference Time 0.396148 0.00699
Unrolling Factor Assignment 0.000001 0.000001
Total Time (s) 0.396313 0.007155
5.1.4. ML Training and Cross Validation
The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test. This ensures no data leakage occurs. Table 4 showcases the classification performance of ACPOLUModel with its top-1 accuracy. As can be seen, ACPOLUModel is able to achieve a geometric mean of 80% accuracy with a standard deviation of 7% by means of its top-1 predictions. In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower. Our investigation reveals the reason lies in the imbalanced training set for the specific cases and it can be alleviated to some extent by employing a number of preprocessing and data augmentation techniques (Wei and Zou, 2019). We leave further investigation to our future works as they are outside the scope of this paper.
Table 4. ACPOLUModel Top-1 Accuracy on Polybench
Benchmark Top-1 Accuracy Benchmark Top-1 Accuracy
2mm 88.411 % gesummv 80 %
3mm 84.254 % gramschmidt 75.466 %
adi 76.543 % lu 85.971 %
atax 57.5 % ludcmp 85.938 %
bicg 63.295 % mvt 75.326 %
cholesky 86.414 % nussinov 79.734 %
correlation 84.091 % seidel-2d 80.909 %
covariance 86.42 % symm 79.797 %
deriche 82.481 % syr2k 84.858 %
doitgen 75.487 % syrk 82.84 %
floyd-warshall 82.864 % trisolv 69.225 %
gemver 86.547 % trmm 90.333 %
Geomean: 80.196%
Establishing a high standalone top-1 accuracy on unseen data is not the sole performance metric to evaluate a model and thus, we deployed the trained model using ACPO infrastructure on LLVM to measure the actual benefits of our ML model. Table 5 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012). We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall. Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
Table 5. ACPOLUModel Speedup on Polybench
Benchmark   Time O3 Var O3 Size O3   Time ACPO Var ACPO Size ACPO Speedup Size Bloat
correlation   2.623 0.29 72728   2.739 0.984 72768 0.957 1.005
covariance   2.743 0.81 72648   2.734 0.927 72728 1.003 1.001
gramschmidt   3.98 0.94 72752   4.371 0.972 72752 0.910 1
cholesky   2.477 0.03 72752   2.248 0.027 72832 1.102 1.001
atax   0.739 0.26 72720   0.749 0.169 72696 0.985 0.999
2mm   3.478 0.97 72608   3.533 0.385 138216 0.984 1.903
3mm   5.182 0.078 72608   5.267 0.752 138248 0.983 1.9040
bicg   1.371 0.025 72664   1.343 0.85 72664 1.020 1
doitgen   0.897 0.044 72680   0.887 0.015 72688 1.011 1.0001
mvt   2.050 0.077 72608   1.959 0.085 72608 1.046 1
gesummv   1.499 0.003 72616   1.311 0.026 72616 1.143 1
gemver   2.699 0.15 72616   2.816 0.151 72648 0.958 1.0004
symm   2.728 0.33 72648   2.572 0.140 72680 1.060 1.0004
syr2k   3.150 0.91 72608   3.176 0.667 72680 0.991 1.0009
syrk   1.500 0.972 72608   1.496 0.879 72648 1.002 1.0005
trmm   1.378 0.92 72648   1.456 0.930 72680 0.946 1.0004
ludcmp   5.040 0.83 72720   4.812 0.6637 72784 1.047 1.0008
lu   5.186 0.31 72696   5.299 0.9422 72816 0.978 1.001
trisolv   1.146 0.013 72648   0.937 0.0171 72648 1.223 1
adi   18.70 0.017 72664   18.478 0.0012 72680 1.012 1.0002
seidel-2d   32.346 0.08 72648   32.318 0.005 72688 1.0008 1.0005
deriche   0.271 0.58 72648   0.251 0.978 72784 1.078 1.0018
floyd-Warshall   28.303 0.06 72624   18.946 0.0217 72656 1.493 1.0004
nussinov   3.839 0.08 72696   3.277 0.12 203808 1.171 2.803
Geomean   - 0.15 -   - 0.237 - 1.040725 1.10193
Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts. ACPO’s prediction provides for a full unroll of a couple of loops including a multiple runtime unrolling of hot loops in the kernel which matches the optimal result our autotuner found after an exhaustive search. O3 on the other hand, only unrolled two loops by their available trip count of 3 and didn’t unroll the remaining hot loops. and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops. Additionally, inspecting our training data generated by means of autotuning, we noticed there was an upper-bound of around 5-6% across Polybench by leveraging the optimal configuration for Loop Unrolling on Aarch64 and thus, achieving an average 4% (shown on Table 5), showcases the performance of the ACPOLUModel on the unseen application of Polybench in a cross-validation mode.
Table 6. ACPOLUModel Generalization Performance
Benchmarks   O3   ACPO   SpUp
Time Var Br Miss Size   Time Var Br Miss Size
[…]
5.1.5. ACPOLUModel Performance Generalization
In Section 5.1.4 we presented the cross-validation results on Polybench, where each application was used as a hold-out set (unseen test case) to evaluate the efficacy of the ACPOLUModel when trained with the remaining Polybench applications. Although the cross-validation results on Polybench show a promising trend for ACPOLUModel, we would still want our model to generalize well across a variety of completely unseen workloads. To this end, we choose a number of applications from Coral-2 benchmarks (Coral-2, 2017), Coremark (Gal-On and Levy, 2012), and Graph 500 (Murphy et al., 2010) to deploy ACPOLUModel on. We aim at selecting a diverse set of standard workloads covering a wide span of large-scale multi-grid solvers to the Embedded domain. Loop unrolling has a complex relation with other subsequent passes in the O3 pipeline, e.g., vectorization, etc, and thus, it is important for an ML model to be able to deduce the kind of correlation between the features derived by the compiler and the target performance metric. This time, we leverage all Polybench data in our training set and build an ACPOLUModel to deploy on the aforementioned benchmarks as test cases.
Table 6 demonstrates our result for which we gain a 5.4% speedup over CoreMark, an average 3% including an 18% speedup over AMG with Coral-2, and a 1.3% speedup on Graph500 SSSP. There are 3 instances, shown in bold under the Var column and by an *** in the Speedup column, that we were not able to contain the variance within 1% by means of multiple try-outs, and thus, we cannot reliably deduce a conclusion for them. Analyzing the obtained results, we notice that the majority of the performance comes from deciding the correct unrolling factor for the hot loops. For instance, Coral-2’s AMG runtime unrolling of a hot loop located at hypre_CSRMatrix class led to an increase in performance by almost 15% which O3 failed to unroll.
5.2. ACPO Function Inlining Model
Similar to Loop Unrolling, Function inlining pass has been optimized by means of adding the necessary code to connect the interface with the ML side but instead of rebuilding an ML model from scratch, we decided to leverage an existing model to showcase the ease of use and the integration capabilities of ACPO when paired with an existing ML model. For this task, we use MLGOPerf (Ashouri et al., 2022) model.
5.2.1. Function Inlining Model
MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm. The model receives as input, 13 function features including block frequency, callsite height, caller and callee block counts and users, etc., and predicts whether or not a callsite is to be inlined. Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology. The model has extended MLGO (Trofin et al., 2021)’s code size reduction to optimize performance instead. As shown in Listing 4, getACPOAdvice() initiates the call to collect necessary features and finally returns the Boolean value of the decision of whether or not to inline a callsite.
5.2.2. Function Inlining Generalization Results
Table 8 showcases the experimental results of the deployed ACPOFIModel when tested with CBench. As expected, the results exhibit similar behavior to MLGOPerf’s experimental results since both are leveraging the same model, albeit differently. For instance, we have gained around 27% on security_sha and inspecting the decisions made by ACPOFIModel, we notice that LLVM’s O3 did not inline the two hot callsites namely, sha_update() and sha_final() whereas in ACPO, those two are inlined. A geometric mean showcases around 2% increase in performance by incurring around 15% code size bloat on Cbench.
5.3. Combined Results
Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench. The combined results have both ACPO’s optimized loop unrolling and function inlining when used in LLVM’s O3 pass pipeline. An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).
Table 7. ACPO Combined Function Inlining & Loop Unrolling Performance (Polybench)
Benchmark Time O3 Var O3 Size O3 Time ACPO Var ACPO Size_ACPO Size Bloat Speedup
2mm 3.472 0.120 72608 3.530 0.023 138144 1.903 0.984
3mm 5.186 0.426 72608 5.311 0.309 203680 2.805 0.976
atax 0.738 0.101 72720 0.748 0.034 72696 1.000 0.987
bicg 1.374 0.081 72664 1.345 0.254 72664 1.000 1.022
cholesky 2.478 0.022 72752 2.330 0.027 72776 1.000 1.063
correlation 2.625 5.276 72728 2.619 0.539 138240 1.901 1.002
covariance 2.743 0.085 72648 2.719 0.988 138184 1.902 1.009
deriche 0.273 0.788 72648 0.245 0.088 138152 1.902 1.115
doitgen 0.897 0.007 72680 0.886 0.015 138192 1.901 1.013
durbin 1.359 0.244 72696 1.395 0.270 72696 1.000 0.974
floyd-warshall 28.314 0.034 72624 18.946 0.011 138160 1.902 1.494
gemver 2.702 0.154 72616 2.772 0.213 72616 1.000 0.975
gesummv 1.503 0.039 72616 1.511 0.019 72616 1.000 0.995
gramschmidt 3.990 0.545 72752 3.988 0.787 138256 1.900 1.000
lu 5.184 0.427 72696 5.207 0.914 72720 1.000 0.996
ludcmp 5.083 0.255 72720 4.767 0.283 72720 1.000 1.066
mvt 2.053 0.042 72608 2.027 0.040 72608 1.000 1.013
nussinov 3.840 0.089 72696 3.270 0.067 72696 1.000 1.174
symm 2.728 0.046 72648 2.544 0.676 138184 1.902 1.072
syr2k 3.136 0.538 72608 3.171 0.971 138144 1.903 0.989
syrk 1.504 0.382 72608 1.452 0.765 138144 1.903 1.036
trisolv 1.149 0.028 72648 0.933 0.098 72648 1.000 1.232
trmm 1.374 0.496 72648 1.409 0.288 72648 1.000 0.975
Geomean - - - - - - 1.383 1.045
Table 8. ACPO Combined Function Inlining & Loop Unrolling Performance (Cbench)
Name ACPO-FI ACPO-LU ACPO Combined
Size Bloat Speedup Size Bloat Speedup Size Bloat Speedup
automotive_bitcount 1.000 0.999 1.001 1.003 1 1.001
automotive_qsort1 1.000 0.994 1.000 0.982 1.91 0.989
automotive_susan_c 1.000 0.999 3.685 1.035 19.797 1.066
automotive_susan_e 1.000 1.023 3.685 1.019 19.797 1.026
automotive_susan_s 1.000 0.980 3.685 1.038 19.797 0.965
bzip2d 1.441 0.991 2.224 1.010 3.541 1.014
bzip2e 1.441 1.007 2.350 1.001 3.546 1.015
consumer_jpeg_c 1.288 1.017 5.385 0.986 5.086 1.027
consumer_jpeg_d 1.289 0.923 5.400 1.073 4.807 1.026
consumer_lame 1.437 1.027 4.501 1.030 4.729 1.034
consumer_mad 1.217 1.071 4.321 0.990 3.193 0.968
consumer_tiff2bw 1.180 1.002 4.502 0.943 2.471 0.987
consumer_tiff2rgba 1.180 1.250 4.681 1.036 2.657 1.012
consumer_tiffdither 1.180 1.046 1.213 0.967 2.655 1.002
consumer_tiffmedian 1.180 1.017 1.000 0.970 3.021 0.943
network_dijkstra 1.000 0.973 1.000 0.990 1 1.037
network_patricia 1.000 0.970 1.000 0.974 1 1.004
office_ghostscript 1.275 1.000 1.000 1.000 1.12 1.001
office_ispell 0.996 1.000 1.000 1.000 1 1
office_rsynth 1.000 1.000 2.024 1.010 2.024 1.011
office_stringsearch1 1.000 1.006 1.000 1.000 1 1
security_blowfish_d 1.000 1.002 1.000 0.978 1 0.965
security_blowfish_e 1.000 1.003 1.000 1.015 1 0.981
security_pgp_d 1.861 0.993 1.000 0.991 13.144 1
security_pgp_e 1.861 1.017 1.000 1.006 13.144 1.006
security_rijndael_d 1.000 0.990 3.823 0.966 1 0.966
security_rijndael_e 1.000 1.001 3.823 0.974 1 0.995
security_sha 0.999 1.270 1.000 1.174 0.999 1.501
telecom_adpcm_c 1.000 1.000 1.000 1.008 1 1.038
telecom_adpcm_d 1.000 0.997 1.000 1.132 1 1.138
telecom_CRC32 1.000 1.094 1.000 1.042 1 1.112
telecom_gsm 1.810 1.076 1.815 1.15 1.81 1.117
Geomean 1.152 1.021 1.818 1.01 2.442 1.024
The results are shown in Tables 7 and 8. We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively. LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Module and Function level for Inline and Loop level for Loop Unroll, is outside the scope of this work. Instead, we showcase the overall benefits of each benchmark. For instance, Table 8 showcases the performance gains using ACPO-FI, ACPO-LU, and the combined version with separate columns.
1//Security_sha Compilation (ACPO)
2...
3Loop Unroll: F[main] Loop %for.body
4    Loop Size = 10
5--- "Huawei’s ACPOModel is activated" ---
6Registering LU features: 30
7Calling ML IF for inference
8Final unroll type post-legality checks is:
9runtime with the unroll count of: 4
10UNROLLING loop %for.body by 4!
11--------------------------------------------
12Loop Unroll: F[byte_reverse] Loop %for.body
13 Loop Size = "15"
14--- "Huawei’s ACPOModel is activated" ---
15Registering LU features: 30
16Calling ML IF for inference
17Final unroll type post-legality checks is:
18runtime with the unroll count of: 8
19COMPLETELY UNROLLING loop %for.body
20with trip count 8!
21...
22--------------------------------------------
23Loop Unroll:F[sha_transform] Loop %for.body5
24  Loop Size = "20"
25--- "Huawei’s ACPOModel is activated" ---
26Registering LU features: 30
27Calling ML IF for inference
28    partially unrolling with count: 16
29UNROLLING loop %for.body5 by 16
30with a breakout at trip 0!
31--------------------------------------------
32Inlining calls in: sha_update
33--- "Huawei’s ACPOModel is activated" ---
34Registering FI features: 17
35Calling ML IF for inference
36ACPOModel inline prediction: 1
37Updated inlining SCC: (sha_update)
38
39--------------------------------------------
40Loop Unroll: F[sha_update] Loop %while.body
41  Loop Size = "457"
42--- "Huawei’s ACPOModel is activated" ---
43Registering LU features: 30
44Calling ML IF for inference
45runtime with the unroll count of: 16
46...
47--------------------------------------------
48Terminating ML interface
49Flushing page cache, dentries and inodes...
50Done flushing.
51...
52Execution Time: 5.3328
Listing 5: ACPO Decisions
1//Security_sha Compilation (LLVM’s O3)
2...
3Loop Unroll: F[main] Loop %for.body
4    Loop Size = 10
5 will not try to unroll loop with runtime trip
6 count -unroll-runtime not given
7
8
9
10
11--------------------------------------------
12Loop Unroll: F[byte_reverse] Loop %for.body
13 Loop Size = "11"
14
15
16
17
18
19COMPLETELY UNROLLING loop %for.body
20with trip count 8!
21...
22--------------------------------------------
23Loop Unroll:F[sha_transform] Loop %for.body5
24  Loop Size = 15
25  Trip Count = 64
26  will not try to unroll partially because
27  -unroll-allow-partial not given
28  "Refer to the explanations in Section 5.3"
29
30
31--------------------------------------------
32Inliner visiting SCC: sha_update: 2 call sites.
33    Inlining (cost=285, threshold=375), Call:
34    call fastcc void @byte_reverse(i64* ...)
35    NOT Inlining (cost=625, threshold=625):
36    call fastcc void @sha_transform(%struct...)
37    NOT Inlining (cost=625, threshold=625):
38    call fastcc void @sha_transform(%struct...)
39--------------------------------------------
40Loop Unroll: F[sha_update] Loop %while.body
41  Loop Size = 73
42  will not try to unroll loop with runtime trip
43 count -unroll-runtime not given
44}
45
46...
47--------------------------------------------
48
49Flushing page cache, dentries and inodes...
50Done flushing.
51...
52Execution Time: 8.012
Listing 6: O3 Decisions
One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 8) using ACPO’s Inliner, now with 50% performance gain. There are a series of decisions made by ACPO models that we highlight an excerpt under Listing 5. Additionally, each corresponding line at Listing 6 shows the decisions made by LLVM’s O3. Note that, partial and runtime unrolling are both disabled by default at LLVM O3 and thus, F[sha_transform] Loop %for.body5 is not unrolled in Listing 6. To be abundantly fair, we measured the performance of O3 when these two flags were forced enabled as well and we noticed that the decision to unroll the loop was changed to 8 as opposed to ACPOLUModel’s prediction of 16 shown in Listing 5. However, evaluations showed that still ACPO was outperforming O3 by a margin of 7% on unrolling this particular benchmark. Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time. Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism as it is less than or equal to the calculated cost. This example showcases the superiority of using 30/13 well-representative loop/function features we designed in training ACPOLUModels to deduce the kind of complex correlation between a code segment and the value of unroll count and the Boolean decision of inlining a callsite to optimize the performance.
6. Discussion
ACPO is the first scalable framework for bringing ML into compilers. It provides the ability to reproduce models on different architecture very fast and the methodology lays out the steps to regenerate data, retrain a new model, and easily deploy with LLVM. Compiler autotuning and optimization space have benefited from ML for almost two decades. As we pointed out earlier, there are many recent works leveraging ML, however, the goal of ACPO is to provide a novel flexible framework for the LLVM community to leverage ML on any optimization passes of interest with minimal efforts. Nevertheless, there are a number of challenges and shortcomings we would like to mention here.
Integration and dependencies
ACPO attempts to alleviate the challenging task of supply chain management and the use of licensed tools by separating the ML framework from the compiler. We showed both the integration of PyTorch and TensorFlow among other moving objects, i.e., Autotuner, into our framework, however, upstreaming would remain challenging and requires community efforts to shape the next generations of ML-enabled compilers. We also provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers. As mentioned in Section 3.6, ACPO uses Python named pipes to perform interprocess communication. Although this allows for greater flexibility, it adds a slight overhead to the compilation time (Shown in Table 3). Developers can switch between these modes when building the compiler. The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
Multi-objective optimization
ACPO optimizes the performance of a code region with intelligent unrolling decisions. However, in a number of cases, we experimentally observe that the emitted code size is also increased. This is a(n) (in)direct result of Inliner and Loop Unroll which increase the number of instructions and thus, the code size. Although this is outside the scope of the current work, an enhanced version may also benefit from taking into account both of the objectives in its exploration strategy by identifying the Pareto optimality (Hoste and Eeckhout, 2008; Ashouri et al., 2013; Moreau et al., 2018; Silva et al., 2021).
7. Conclusion
In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM. We demonstrated that using ACPO framework, we are able to enable Loop Unroll pass to benefit from a seamless integration of an ML model, i.e., ACPOLUModel, for which we experimentally achieve on average a 4%, 3%, 5.4%, and 0.2% performance speedup over LLVM’s O3 on Polybench, Coral-2, CoreMark, and Graph-500, respectively. Additionally, we showcased a scenario where we leverage an existing ML model to optimize Function Inlining pass and similar to the original model’s performance, we experimentally obtained 2% speedup wrt LLVM’s O3 when deployed on CBench and we showed that by combining both models, the benefits increase up to 4.5% and 2.4% on Polybench and Cbench, respectively. Future works will be focused on extending ACPO framework by enabling a number of other optimization passes inside LLVM’s pipeline and leveraging innovative software characterization methods to be employed.
Appendix A Appendix —  ACPO Headers
1class ACPOModel {
2public:
3  ACPOModel(...);
4  ~ACPOModel();
5  void setMLIF(std::shared_ptr<ACPOMLInterface> ML);
6  std::shared_ptr<ACPOMLInterface> getMLIF();
7  void addRequiredResultField(std::string Name, Type::TypeID &ID);
8  std::unique_ptr<ACPOAdvice> getAdvice();
9  void addFeature(int64_t ID, Constant *Val);
10  void setCustomFeatures(const std::vector<std::pair<std::string, Constant *>> &FeatureValues);
11
12protected:
13  virtual void prepareModelInput();
14  virtual bool runModel(std::unique_ptr<ACPOAdvice> &Result);
15};
Listing 7: ACPOModel Header
1class ACPOMLInterface {
2public:
3  ACPOMLInterface();
4  virtual ~ACPOMLInterface();
5  virtual bool loadModel(std::string ModelSpecFile);
6  virtual bool freeModel(std::string ModelName);
7  virtual bool registerModel(std::string ModelName, int NumFeatures, int NumOutputs);
8  virtual bool registerFeature(std::string ModelName, std::string FeatureName, int Index);
9  virtual bool registerOutput(std::string ModelName, std::string OutputName, std::string OutputType);
10  virtual bool setCustomFeature(std::string ModelName, uint64_t FeatureID, int FeatureValue);
11  ...
12  virtual bool setCustomFeature(std::string ModelName, std::string FeatureName, bool FeatureValue);
13  virtual bool setCustomFeatures(std::string ModelName, const std::vector<std::pair<uint64_t, std::string>> &FeatureValues);
14  virtual bool setCustomFeatures(std::string ModelName, const std::vector<std::pair<std::string, std::string>> &FeatureValues);
15  ...
16  virtual bool runModel(std::string ModelName);
17  virtual std::string getOutputType(std::string ModelName, std::string OutputName);
18  virtual int getModelResultI(std::string OutputName);
19  ...
20  virtual bool getModelResultB(std::string OutputName);
21  ...
22  virtual int getStatus();
23  virtual bool closeMLInterface();
24};
Listing 8: ACPOMLInterface Header

### 11 Mar 2024 v2

@@
 The key to performance optimization of a program is to decide correctly when a certain transformation should be applied by a compiler.
-Traditionally, such profitability decisions are made by hand-coded algorithms tuned for a very small number of benchmarks, usually requiring a great deal of effort to be retuned when the benchmark suite changes.
 This is an ideal opportunity to apply machine-learning models to speed up the tuning process; while this realization has been around since the late 90s, only recent advancements in ML enabled a practical application of ML to compilers as an end-to-end framework.
-Even so, seamless integration of ML into the compiler would require constant rebuilding of the compiler when models are updated.
 This paper presents ACPO: AI-Enabled Compiler-driven Program Optimization; a novel framework to provide LLVM with simple and comprehensive tools to benefit from employing ML models for different optimization passes.
 We first showcase the high-level view, class hierarchy, and functionalities of ACPO and subsequently, demonstrate a couple of use cases of ACPO by ML-enabling the Loop Unroll and Function Inlining passes and describe how ACPO can be leveraged to optimize other passes.
-Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% and 3%, 5.4%, 0.2% compared to LLVM’s O3 optimization when deployed on Polybench, Coral-2, CoreMark, and Graph-500, respectively.
+Experimental results reveal that ACPO model for Loop Unroll is able to gain on average 4% compared to LLVM’s O3 optimization when deployed on Polybench.
 Furthermore, by adding the Inliner model as well, ACPO is able to provide up to 4.5% and 2.4% on Polybench and Cbench compared with LLVM’s O3 optimization, respectively.
 Compilers, Optimization Framework, Loop Unrolling, Inlining, Deep Learning, LLVM
-copyright: nonejournal: tacoccs: Software and its engineering Compilersccs: Computing methodologies Machine learning{textblock}
-13(1.6,1) Under revision at ACM Transactions on Architecture and Code Optimization (TACO)
-Received 26 June 2023; Revised 20 Nov 2023
-LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
+copyright: noneccs: Software and its engineering Compilersccs: Computing methodologies Machine learning{textblock}
+30(2,0.5) LLVM Dev 2023 Presentation: https://youtu.be/jzT-tar_X0U
 1. Introduction
@@
 Leveraging machine learning (ML) for compilers has become an extensively researched topic in the past years.
-In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023).
+In many instances, the authors showed experimental results that outperformed LLVM’s O3 (Agakov et al., 2006; Chen et al., 2012; Park et al., 2013; Ashouri et al., 2016, 2017; Cummins et al., 2017; Georgiou et al., 2018; Haj-Ali et al., 2020; Ashouri et al., 2022) or Deep Neural Networks (DNN) kernel schedulers (Patabandi et al., 2021; Patabandi and Hall, 2023) and lately, using Large Language Models (LLMs) (Cummins et al., 2023; Gong et al., 2024).
 Machine learning approaches normally rely on data generated by an autotuner (Ansel et al., 2014; Kalyan et al., 2019), which helps with exploring an optimization space formed by the number of parameters in the decision set.
@@
 In recent years, there have been considerable advancements in the space, especially in the graph-based program representation techniques (Guo et al., 2020; Cummins et al., 2021a) and token-based identifiers (Wang et al., 2021), which are outside the scope of this work.
-Recently, there have been a few works providing an end-to-end approach to construct ML-based heuristics for optimizing the decision making process of the transformation passes (Cummins et al., 2017; Das et al., 2020; Rotem and Cummins, 2021; Trofin et al., 2021; Cummins et al., 2021b; Mannarswamy and Das, 2022; Ashouri et al., 2022).
-Among these, only a subset provides an ML-guided optimization approach (Cummins et al., 2021b; Rotem and Cummins, 2021; Trofin et al., 2021; Ashouri et al., 2022) baked into a single or multiple passes to readily streamline the optimization process inside the pass functionality.
-The recent advancements in this area beg the question of how scalable, user-friendly, and comprehensive a framework should be, to provide seamless integration with existing LLVM optimization passes, while maintaining a clean separation of concerns between the ML and the LLVM layers.
 The vision for ACPO is to enable the use of Machine Learning in a compiler, in a way that is modular, extendable, and easy to use.
@@
 Section 2 discusses the state of the art.
-Section 3 gives details on our proposed framework and how the different layers of abstraction communicate with each other.
-Under Section 4.1, we showcase how Loop Unroll pass can benefit by instantiating ACPOLUModel and, subsequently, Section 4.2 does so for Function Inlining pass.
+Section 3 gives details on our proposed framework.
+Under Sections 4.1 and 4.2, we showcase how Loop Unroll and Function Inlining passes can benefit by instantiating ACPOLUModel.
 Section 5 showcases the experimental results including the individual and the combined results of deploying both models and Finally, in Section 6 we discuss challenges and propose directions for future work.
 2. Related Work
-A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018a) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics.
+A recent survey on the compiler autotuning and optimization (Ashouri et al., 2018) reveals that there have been a vast number of approaches to employ ML techniques for automatically constructing optimization heuristics.
 These approaches mainly focused on two major problems of compiler optimization: (1) the optimization selection (Bodin et al., 1998; Cooper et al., 2002; Agakov et al., 2006; Fursin and Temam, 2009; Park et al., 2013; Ashouri et al., 2016) and (2) the phase ordering problem (Kulkarni and Cavazos, 2012; Ashouri et al., 2017; Nobre et al., 2018; Huang et al., 2019; Mammadli et al., 2020; Cummins et al., 2021b; Wang et al., 2022a).
@@
 Additionally, our objective is to improve the performance as opposed to reducing the code size of a binary.
-Rotem and Cummins (Rotem and Cummins, 2021) propose a Decision Trees (DS)-based framework to infer branch probabilities for use in providing profile-guided optimizations (PGO).
-The authors use hand-crafted features for the characterization of code regions and show they are able to reap the benefits of PGO without actually running the program with instrumentation to collect profile data and recompiling.
-This work is relevant to ours since the approach builds into LLVM and once trained, it can provide PGO certain profiling information without actually running PGO.
 Trofin et al.
@@
 Additionally, the authors have extended the code base to provide optimization for the register allocation problem.
-Our previous work, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g.
+Later, MLGOPerf (Ashouri et al., 2022), extended MLGO to optimize inlining for performance as opposed to code size and showed efficacy on a number of benchmarks, e.g.
 SPEC CPU 2006 and Cbench.
@@
 Figure 2. ACPO Flow
-3.1. Autotuner
-An autotuner is a software tool that derives decisions for an optimization scenario (minimization or maximization), by intelligently exploring the design space of compiler optimization parameters (Palermo et al., 2005; Ansel et al., 2014; Ashouri et al., 2018a, b).
-The parameters can be either coarse- or fine-grained.
+Autotuner
+The autotuner parameters can be either coarse- or fine-grained.
 The former includes the selection and the phase-ordering of different optimization passes as a black box, as well as optimization parameters that apply to the whole translation unit, while the latter facilitates the tuning of optimization parameters for individual code regions of a program, e.g.
@@
 To generate data for training purposes we rely on the autotuner.
-The autotuning process allows us to override compiler heuristics with an iterative search for the optimal optimization decisions for a given application.
-By performing the search on our benchmark programs and recording each step of the search, we can obtain information about the impact of each decision on program performance, and thus collect sufficient information to be able to train the eventual model with a reasonable chance of success.
 To do this efficiently, we enabled what we refer to as fine-grained autotuning to control each decision of a given transformation pass independently of one another.
@@
 With the help of fine-grained autotuning, we are able to specify which pass we wish to tune and apply this kind of tuning to a set of benchmarks.
-The first step in the autotuning flow is to establish a search space over which the tuning occurs.
-This comprises both the focus on a pass as well as code regions on which the specific transformation can act.
-In loop unrolling for instance, the applicable code regions would be loops, not single instructions or whole functions.
-Once the search space is established, the autotuner simply proposes solutions for each code region and then requests the compiler to apply the desired optimization on the given code region at the appropriate time.
-This is crucial because we do not want to affect the compiler pipeline itself; instead, we want to generate a wide array of solutions at a specific point in the compilation process and then observe the impact of those decisions.
-That is key to creating a viable model.
-At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions.
-Initially, the pass proceeds normally, starting with legality checks and then following through to profitability analysis, which is when the ML model will be invoked.
-The compiler instantiates an ACPOModel and gathers features pertaining to the code region being optimized.
-This is important to ensure that the model being trained is provided with input features based on what they will be at deployment time when the compiler uses this ML-enabled pass.
-Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation.
+At the compilation stage, with the instructions from the autotuner, the compiler proceeds to optimize an application and eventually reaches the ML-enabled pass for which the autotuner is currently providing decisions.Note that ACPO supports leveraging different scopes of features as long as it is supported by the compiler and it is up to the developer to decide which ACPO pass might benefit from which scope during the process of design and implementation.
 The features are passed with std::vector of pairs to the ML side.
 This process can also be scaled to an arbitrary number of features and feature types.
-The feature set is then saved for later use to allow the features, the decision specified by the autotuner, and the final impact of the decision to be correlated.
 The transformation decision as prescribed by the autotuner is applied to the code region (e.g.
@@
 Once the compilation is completed, the autotuner runs the program and collects a measurement for performance feedback (the simplest metric being the execution time of the program).
-The measurement is used as an assessment of decision impact, so the decisions that lead to a positive impact can be distinguished from the ones that lead to a negative impact on performance.
-Generally, we use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation.
-This process is repeated many times to gather as much data per benchmark as possible to achieve beneficial results on most benchmarks.
-The resulting trial data is then processed and used as input to the model training.
+We use the O3 compilation as our baseline and measure impact in terms of speedup relative to the O3 compilation.
 3.2. ACPO Inference
@@
 In this work we focus on improving decision making within a pass using ML, so changes to the pipeline ordering and optimization selection of passes are orthogonal and outside the scope of the work.
-One of the key things we added to the compiler is the ability to invoke a model and run inference given a set of features.
-To facilitate this, ML-enabled passes in the compiler have been given access to ACPOModel class, from which a pass-specific model instance can be derived.
 3.4. ACPOModel
 The ACPOModel class is responsible for storing model-specific information and leveraging ML APIs to performing inference.
-The header file snippets of ACPOModel can be found in the Appendix section under Listing 7.
+The header file code snippet is left out of this version for brevity, but it contains the same functionality described in figures 1 and 2.
 This class is a simple abstraction of how the model is incorporated into the compiler flow.
@@
 The reason for this is to enable derivation of this class to represent a new custom model and allow the developer to alter the flow in a way consistent with how they envision the use of the model and how they wish to handle errors.
-3.6. ML APIs
+3.5. ML APIs & Framework
 An important part of the ACPOModel class is the reference to the ACPOMLInterface class.
 This class is the basis for the ML API in our framework and separates the actual ML model details by abstracting the underlying ML framework away from ACPOModel class, and the compiler.
-The snippet for the ACPOMLInterface header is shown under the Appendix section in Listing 8.
+The code snippet is left out of this version for brevity, but it contains the same functionality described in figure 2.
 The ACPOMLInterface class is virtual and it is meant to be implemented depending on how your particular ML framework needs are accessed.
@@
 os.mkfifo().
-The ML framework in our scenario is a separate python process that has a user-visible interface through the named pipes.
-It receives commands from ACPOMLInterface-derived class on the compiler side, processes the requests, and responds with requested information or status.
 This approach allows us to separate the compiler from the ML framework completely, allowing the model implementation to be on the Python side and for it to be changed without modifying the compiler on a regular basis.
 While at the moment the Python side of the ACPO framework is very lightweight, its separation from the compiler permits easy build-up of tools and utilities to support more models and model types without compiler changes.
-We envision this will grow to comprise a toolkit with many different models, algorithms, and program representation mechanisms at the developer’s disposal.
 Later in the experimental evaluation section, we showcase the overhead of different component of ACPO in Table 1.
-3.7. ML Framework
-In this work, to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment.
-Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects.
+Note that users can leverage any ML framework of their choice as long as the input/outputs match what the interface expects and in this work to showcase the versatility of ACPO, we used two different ML frameworks, i.e., PyTorch -- for model design and training and, TensorFlow -- for model deployment.
 4. ACPO Models
@@
 4.1. ACPO Loop Unrolling Model
-Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018a).
-There have been multiple works trying to automatically construct heuristics to find the optimal unrolling factor (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018) and research has shown identifying the optimal unroll count is a fine line between the increase in speed versus the binary code size by which the increase in cache misses can be counterproductive.
-Therefore, a number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.
+Identifying the optimal unrolling value by which a loop needs to be unrolled is known to be a hard problem with a non-polynomial complexity (Sarkar, 2000; Ashouri et al., 2018).
+There have been multiple works trying to automatically construct heuristics to find the optimal unrolling count (Sarkar, 2000; Lokuciejewski and Marwedel, 2009; Murthy et al., 2010; Zacharopoulos et al., 2018).
+A number of factors need to be examined by the decision maker to ensure the decrease in the number of control instructions, i.e., branch misses, can outweigh the increase in the number of overall instructions and eventually, the code size bloat, i.e., higher cache misses.
 1class ACPOLUModel : public ACPOModel {
@@
 A new ACPOModel class, namely ACPOLUModel can be instantiated as shown under Listing 1.
-We start by providing the necessary arguments for the constructor and initializing the private members.
-Subsequently, we run a few checks to make sure the steps mentioned prior under Section 3.6 are performed correctly.
-That leads us to the main inference method which is getAdviceML().
-This method first retrieves a reference to the ML interface which is used to communicate with the ML framework.
+The textttgetAdviceML() method first retrieves a reference to the ML interface which is used to communicate with the ML framework.
 It then loads a model used for inference, which causes the framework to load the appropriate model as specified in the model.acpo file.
-If the model has already been loaded by the framework, then this command completes instantly, as loading the same model many times serves no purpose.
 The method then sets the inputs to the model and invokes the runModel method to begin inference.
@@
 However, if for any reason inference fails, the result of runModel will be false, indicating that the process failed.
-This mechanism allows the compiler to handle this case as the developer chooses.
-In our case, we forced the compiler to terminate with an error to ensure our ML model was the only mechanism for determining profitability.
-However, developers are welcome to provide fall-back mechanisms as they see fit.
-To leverage ML in unrolling, we invoked the ACPOLUModel after the loop unrolling legality checks had been made.
-Also, we respected the high-priority overrides.
-The highest level is set by the explicit use of unroll-count by the user, followed by the usage of #pragma clang loop unroll_count(N).
-We invoke ACPOLUModel object subsequent to these two priorities and provide the unrolling information to the pass as shown in Listing 2.
 1  if (EnableACPOLU.getNumOccurrences() > 0 && LAA) {
@@
 Function inlining is one of the fundamental compiler optimizations used by many state-of-the-art compilers.
-In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018a; Theodoridis et al., 2022).
+In recent years, several attempts have been made to leverage ML on optimizing the pass for both code size reduction (Trofin et al., 2021; Theodoridis et al., 2022) and performance (Ashouri et al., 2022) and it was shown the optimal function inlining problem has an exponential increase in optimization space (Ashouri et al., 2018; Theodoridis et al., 2022).
 A new ACPOModel class, namely, ACPOFIModel can be instantiated as shown under Listing 3 .
@@
 Listing 3: ACPOFIModel Instantiation
-To leverage ML in function inlining, we invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made.
+We invoked the ACPOFIModel by adding getACPOAdvice() after the legality checks have been made.
 Also, we respected the high-priority overrides.
@@
 5. Experimental Evaluation
-In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of (1) standalone prediction accuracy, (2) cross-validation method using Polybench, and (3) the generalization evaluating of the model using a number of other standard benchmarks.
-Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
-5.1. ACPO Loop Unrolling Model
+In this Section, we provide details on our experimental campaign to showcase the efficacy of both ACPOLUModel by means of cross-validation method on Polybench Subsequently, we deploy and showcase the results we reproduced with ACPOFIModel.
+ACPO Loop Unrolling Model
 As mentioned earlier, Loop Unroll pass has been optimized from scratch by means of adding the necessary code to connect the interface with the ML side.
@@
 The details of the design and testing are provided as follows.
-5.1.1. ACPOLUModel Architecture
-Table 1 presents the architecture of ACPOLUModel.
-It is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor.
+ACPOLUModel is a 5-layer Neural Net with ReLu as the activation layer and a final Softmax layer to perform the classification task to identify the optimal unrolling factor.
 i.e., UP.Count values, given a loop.
@@
 As mentioned earlier in Section 3.5, users can use any ML infrastructure of choice when training and deploying models as long as the input and output between the ML APIs and the compiler side match.
-Table 1. ACPO Loop Unrolling Model Architecture
-Layer No
-Layer (type) Output Shape
-0
-Linear-FC1 [Features, 1, 32]
-Relu-0 [Features, 1, 32]
-1
-Linear-FC1 [-1, 1, 128]
-Relu-1 [-1, 1, 128]
-2
-Linear-FC2 [-1, 1, 256]
-Relu-2 [-1, 1, 256]
-3
-Linear-FC3 [-1, 1, 64]
-Relu-3 [-1, 1, 64]
-4
-Linear-FC4 [-1, 1, Classes]
-Softmax [-1, 1, Classes]
 The input to the model is a set of 30 handcrafted features, derived from the LLVM IR to characterize the state of any given loop being transformed.
@@
 Additionally, the framework supports adding other types of features, i.e., graph-based, tokens, etc., however, in this work, we leverage handcrafted features derived from LLVM IR.
-The features are shown in Table 2.
+The features used are shown in Table 3 and taken from our super class of ACPOFeatures which contains a large set of handcrafted features defined in different scopes, i.e., Module, Function, Call Graph, and Loops, so that users can easily add and register them into their optimization use cases.
 The Pearson (Benesty et al., 2009) correlation analysis between the features and the target label, i.e.
-classes corresponding to unrolling count, is depicted in Figure 3.
+classes corresponding to unrolling count, is also depicted in Figure 3.
 The correlation values are within the range of [-1, +1], and specifically, we aimed at designing features with the highest positive/negative correlation with the target label.
-We did this by using features with at least a $\pm$5correlation, and this can be seen in the final column/row.
-We collect these features prior to the actual loop unrolling transformation routine and leverage setCustomFeatures() method to communicate with the ML side via the MLInterface API.
-Note that feature engineering and software characterization (Fursin et al., 2011; Ding et al., 2019; Cummins et al., 2021a) is a dynamic and fast-paced topic in the space and is outside the scope of this work.
-We leave the usage of other methods for our future work.
-Table 2. ACPOLUModel Features
-No Static Features No Static Features
-1
-PartialOptSizeThreshold
-16
-NumLoadInstPerLoopNest
-2
-AllowRemainder
-17
-NumStoreInstPerLoopNest
-3
-UnrollRemainder
-18
-TotLoopNestInstCount
-4
-AllowExpensiveTripCount
-19
-AvgNumLoadInstPerLoopNest
-5
-Force
-20
-AvgNumStoreInstPerLoopNest
-6
-TripCount
-21
-NumLoadInstPerLoop
-7
-MaxTripCount
-22
-NumStoreInstPerLoop
-8
-Size
-23
-TotLoopInstCount
-9
-InitialIVValueInt
-24
-AvgNumLoadInstPerLoop
-10
-FinalIVValueInt
-25
-AvgNumStoreInstPerLoop
-11
-StepValueInt
-26
-IsInnerMostLoop
-12
-NumPartitions
-27
-IsOuterMostLoop
-13
-IndVarSetSize
-28
-MaxLoopHeight
-14
-AvgStoreSetSize
-29
-TotBlocksPerLoop
-15
-AvgNumInsts
-30
-IsFixedTripCount
+We did this by using features with at least a $\pm$2correlation.
 Figure 3. Features Correlation Analysis
-5.1.2. Data Collection
+Data Collection
 We used the autotuner to collect a total of 10000 different unrolling configurations from Polybench by tuning the unroll count of individual loops (UP.Count), where the set of UP.Count values are $\in\{0,2,\cdots,64\}$.
@@
 Then, the compiler uses the configurations suggested by the autotuner and generates a new executable file.
-We run the emitted executable and provide the execution times to the autotuner as feedback which will be leveraged by the autotuner to generate a new set of configurations based on the feedback.
 This process was repeated for 100 iterations to generate our training data.
@@
 In this work, we define the Loop Unrolling speedup as the following:
-(1) $\footnotesize Loop_{speedup}=\frac{Loop_{runtime(Base)}}{Loop_{runtime(X)}}$
+(1) $\scriptsize Loop_{speedup}=\frac{Loop_{runtime(Base)}}{Loop_{runtime(X)}}$
 where $Loop_{runtime(Base)}$ is the execution time of a module having multiple loops at LLVM O3’s configuration and Loop_runtime(X), the same module having a loop unrolling count configuration of X in the set of all Unrolling configurations generated by our autotuner.
@@
 This way ACPOLUModel only exposes to the set of quality samples and is trained to learn the relation between the given features and the target UP.Count.
-5.1.3. Analysis of Compilation Time Overhead
+Analysis of Compilation Time Overhead
 We measure the compilation-time overhead of utilizing ACPOLUModel with a breakdown of each individual task, as shown in Listing 2, and we present the breakdown in Table 1.
@@
 Total Time (s) 0.396313 0.007155
-5.1.4. ML Training and Cross Validation
+ML Training and Cross Validation
 The training is carried out using leave-one-out cross-validation for which we exclude one application from the training data and use it later as the test.
 This ensures no data leakage occurs.
-Table 4 showcases the classification performance of ACPOLUModel with its top-1 accuracy.
-As can be seen, ACPOLUModel is able to achieve a geometric mean of 80% accuracy with a standard deviation of 7% by means of its top-1 predictions.
+Measured Top-1 accuracy of the ACPOLUModel’s predictions exhibits a geometric mean of 80% accuracy with a standard deviation of 7% on Polybench.
 In certain applications, e.g., trmm, 2mm, etc., the top-1 accuracy is higher than the average, and in a few such as atax, it is lower.
@@
 We leave further investigation to our future works as they are outside the scope of this paper.
-Table 4. ACPOLUModel Top-1 Accuracy on Polybench
-Benchmark Top-1 Accuracy Benchmark Top-1 Accuracy
-2mm 88.411 % gesummv 80 %
-3mm 84.254 % gramschmidt 75.466 %
-adi 76.543 % lu 85.971 %
-atax 57.5 % ludcmp 85.938 %
-bicg 63.295 % mvt 75.326 %
-cholesky 86.414 % nussinov 79.734 %
-correlation 84.091 % seidel-2d 80.909 %
-covariance 86.42 % symm 79.797 %
-deriche 82.481 % syr2k 84.858 %
-doitgen 75.487 % syrk 82.84 %
-floyd-warshall 82.864 % trisolv 69.225 %
-gemver 86.547 % trmm 90.333 %
-Geomean: 80.196%
-Establishing a high standalone top-1 accuracy on unseen data is not the sole performance metric to evaluate a model and thus, we deployed the trained model using ACPO infrastructure on LLVM to measure the actual benefits of our ML model.
-Table 5 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012).
+Subsequent to establishing the high top-1 accuracy on a cross validation mode, we evaluate the performance of the ACPOLUModel by means of deploying it on Polybench and measuring the actual benefits of our ML model.
+Table 2 showcases the ACPOLUModel’s results when deployed on Polybench (Pouchet, 2012) on Aarch64.
 We observe that on average, ACPOLUModel achieved a speedup of 4.07% among the Polybench benchmarks with a best of 49.3% on Floyd-Warshall.
 Additionally, we observe that out of 24 benchmarks, ACPOLUModel can outperform LLVM’s O3 on 15, excluding the close call for syr2k application.
-Table 5. ACPOLUModel Speedup on Polybench
-Benchmark   Time O3 Var O3 Size O3   Time ACPO Var ACPO Size ACPO Speedup Size Bloat
-correlation   2.623 0.29 72728   2.739 0.984 72768 0.957 1.005
-covariance   2.743 0.81 72648   2.734 0.927 72728 1.003 1.001
-gramschmidt   3.98 0.94 72752   4.371 0.972 72752 0.910 1
-cholesky   2.477 0.03 72752   2.248 0.027 72832 1.102 1.001
-atax   0.739 0.26 72720   0.749 0.169 72696 0.985 0.999
-2mm   3.478 0.97 72608   3.533 0.385 138216 0.984 1.903
-3mm   5.182 0.078 72608   5.267 0.752 138248 0.983 1.9040
-bicg   1.371 0.025 72664   1.343 0.85 72664 1.020 1
-doitgen   0.897 0.044 72680   0.887 0.015 72688 1.011 1.0001
-mvt   2.050 0.077 72608   1.959 0.085 72608 1.046 1
-gesummv   1.499 0.003 72616   1.311 0.026 72616 1.143 1
-gemver   2.699 0.15 72616   2.816 0.151 72648 0.958 1.0004
-symm   2.728 0.33 72648   2.572 0.140 72680 1.060 1.0004
-syr2k   3.150 0.91 72608   3.176 0.667 72680 0.991 1.0009
-syrk   1.500 0.972 72608   1.496 0.879 72648 1.002 1.0005
-trmm   1.378 0.92 72648   1.456 0.930 72680 0.946 1.0004
-ludcmp   5.040 0.83 72720   4.812 0.6637 72784 1.047 1.0008
-lu   5.186 0.31 72696   5.299 0.9422 72816 0.978 1.001
-trisolv   1.146 0.013 72648   0.937 0.0171 72648 1.223 1
-adi   18.70 0.017 72664   18.478 0.0012 72680 1.012 1.0002
-seidel-2d   32.346 0.08 72648   32.318 0.005 72688 1.0008 1.0005
-deriche   0.271 0.58 72648   0.251 0.978 72784 1.078 1.0018
-floyd-Warshall   28.303 0.06 72624   18.946 0.0217 72656 1.493 1.0004
-nussinov   3.839 0.08 72696   3.277 0.12 203808 1.171 2.803
-Geomean   - 0.15 -   - 0.237 - 1.040725 1.10193
 Further investigation on our obtained results reveals that Floyd-Warshall application has a seven-loop module and the added loops from polybench.c with two instances of loop unrolling pass under O3 makes it for an opportunity set of 23 loop-nest regions which can be tuned by different unroll counts.
@@
 and thus, we noticed that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
-Additionally, inspecting our training data generated by means of autotuning, we noticed there was an upper-bound of around 5-6% across Polybench by leveraging the optimal configuration for Loop Unrolling on Aarch64 and thus, achieving an average 4% (shown on Table 5), showcases the performance of the ACPOLUModel on the unseen application of Polybench in a cross-validation mode.
-Table 6. ACPOLUModel Generalization Performance
-Benchmarks   O3   ACPO   SpUp
-Time Var Br Miss Size   Time Var Br Miss Size
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-5.1.5. ACPOLUModel Performance Generalization
-In Section 5.1.4 we presented the cross-validation results on Polybench, where each application was used as a hold-out set (unseen test case) to evaluate the efficacy of the ACPOLUModel when trained with the remaining Polybench applications.
-Although the cross-validation results on Polybench show a promising trend for ACPOLUModel, we would still want our model to generalize well across a variety of completely unseen workloads.
-To this end, we choose a number of applications from Coral-2 benchmarks (Coral-2, 2017), Coremark (Gal-On and Levy, 2012), and Graph 500 (Murphy et al., 2010) to deploy ACPOLUModel on.
-We aim at selecting a diverse set of standard workloads covering a wide span of large-scale multi-grid solvers to the Embedded domain.
-Loop unrolling has a complex relation with other subsequent passes in the O3 pipeline, e.g., vectorization, etc, and thus, it is important for an ML model to be able to deduce the kind of correlation between the features derived by the compiler and the target performance metric.
-This time, we leverage all Polybench data in our training set and build an ACPOLUModel to deploy on the aforementioned benchmarks as test cases.
-Table 6 demonstrates our result for which we gain a 5.4% speedup over CoreMark, an average 3% including an 18% speedup over AMG with Coral-2, and a 1.3% speedup on Graph500 SSSP.
-There are 3 instances, shown in bold under the Var column and by an *** in the Speedup column, that we were not able to contain the variance within 1% by means of multiple try-outs, and thus, we cannot reliably deduce a conclusion for them.
-Analyzing the obtained results, we notice that the majority of the performance comes from deciding the correct unrolling factor for the hot loops.
-For instance, Coral-2’s AMG runtime unrolling of a hot loop located at hypre_CSRMatrix class led to an increase in performance by almost 15% which O3 failed to unroll.
 5.1. ACPO Function Inlining Model
@@
 MLGOPerf (Ashouri et al., 2022) is a reinforcement learning model trained with Proximal Policy Optimization (PPO) (Schulman et al., 2017) algorithm.
-The model receives as input, 13 function features including block frequency, callsite height, caller and callee block counts and users, etc., and predicts whether or not a callsite is to be inlined.
+The model receives as input, 13 function features including block frequency, callsite height, block counts, etc., and predicts whether or not a callsite is to be inlined.
 Readers can refer to the Table 2 of the reference paper for the exhaustive list of features used and the training methodology.
@@
 Finally, we deploy both ACPO models in combined mode and experimentally observe the benefits on Polybench and Cbench.
-The combined results have both ACPO’s optimized loop unrolling and function inlining when used in LLVM’s O3 pass pipeline.
 An inspection for O3’s pipeline reveals that function inlining triggers before the first instance of loop unrolling and thus, an enhanced function inlining may benefit the subsequent passes, including loop unrolling (Ashouri et al., 2022).
-Table 7. ACPO Combined Function Inlining & Loop Unrolling Performance (Polybench)
-Benchmark Time O3 Var O3 Size O3 Time ACPO Var ACPO Size_ACPO Size Bloat Speedup
-2mm 3.472 0.120 72608 3.530 0.023 138144 1.903 0.984
-3mm 5.186 0.426 72608 5.311 0.309 203680 2.805 0.976
-atax 0.738 0.101 72720 0.748 0.034 72696 1.000 0.987
-bicg 1.374 0.081 72664 1.345 0.254 72664 1.000 1.022
-cholesky 2.478 0.022 72752 2.330 0.027 72776 1.000 1.063
-correlation 2.625 5.276 72728 2.619 0.539 138240 1.901 1.002
-covariance 2.743 0.085 72648 2.719 0.988 138184 1.902 1.009
-deriche 0.273 0.788 72648 0.245 0.088 138152 1.902 1.115
-doitgen 0.897 0.007 72680 0.886 0.015 138192 1.901 1.013
-durbin 1.359 0.244 72696 1.395 0.270 72696 1.000 0.974
-floyd-warshall 28.314 0.034 72624 18.946 0.011 138160 1.902 1.494
-gemver 2.702 0.154 72616 2.772 0.213 72616 1.000 0.975
-gesummv 1.503 0.039 72616 1.511 0.019 72616 1.000 0.995
-gramschmidt 3.990 0.545 72752 3.988 0.787 138256 1.900 1.000
-lu 5.184 0.427 72696 5.207 0.914 72720 1.000 0.996
-ludcmp 5.083 0.255 72720 4.767 0.283 72720 1.000 1.066
-mvt 2.053 0.042 72608 2.027 0.040 72608 1.000 1.013
-nussinov 3.840 0.089 72696 3.270 0.067 72696 1.000 1.174
-symm 2.728 0.046 72648 2.544 0.676 138184 1.902 1.072
-syr2k 3.136 0.538 72608 3.171 0.971 138144 1.903 0.989
-syrk 1.504 0.382 72608 1.452 0.765 138144 1.903 1.036
-trisolv 1.149 0.028 72648 0.933 0.098 72648 1.000 1.232
-trmm 1.374 0.496 72648 1.409 0.288 72648 1.000 0.975
-Geomean - - - - - - 1.383 1.045
-Table 8. ACPO Combined Function Inlining & Loop Unrolling Performance (Cbench)
-Name ACPO-FI ACPO-LU ACPO Combined
-Size Bloat Speedup Size Bloat Speedup Size Bloat Speedup
-automotive_bitcount 1.000 0.999 1.001 1.003 1 1.001
-automotive_qsort1 1.000 0.994 1.000 0.982 1.91 0.989
-automotive_susan_c 1.000 0.999 3.685 1.035 19.797 1.066
-automotive_susan_e 1.000 1.023 3.685 1.019 19.797 1.026
-automotive_susan_s 1.000 0.980 3.685 1.038 19.797 0.965
-bzip2d 1.441 0.991 2.224 1.010 3.541 1.014
-bzip2e 1.441 1.007 2.350 1.001 3.546 1.015
-consumer_jpeg_c 1.288 1.017 5.385 0.986 5.086 1.027
-consumer_jpeg_d 1.289 0.923 5.400 1.073 4.807 1.026
-consumer_lame 1.437 1.027 4.501 1.030 4.729 1.034
-consumer_mad 1.217 1.071 4.321 0.990 3.193 0.968
-consumer_tiff2bw 1.180 1.002 4.502 0.943 2.471 0.987
-consumer_tiff2rgba 1.180 1.250 4.681 1.036 2.657 1.012
-consumer_tiffdither 1.180 1.046 1.213 0.967 2.655 1.002
-consumer_tiffmedian 1.180 1.017 1.000 0.970 3.021 0.943
-network_dijkstra 1.000 0.973 1.000 0.990 1 1.037
-network_patricia 1.000 0.970 1.000 0.974 1 1.004
-office_ghostscript 1.275 1.000 1.000 1.000 1.12 1.001
-office_ispell 0.996 1.000 1.000 1.000 1 1
-office_rsynth 1.000 1.000 2.024 1.010 2.024 1.011
-office_stringsearch1 1.000 1.006 1.000 1.000 1 1
-security_blowfish_d 1.000 1.002 1.000 0.978 1 0.965
-security_blowfish_e 1.000 1.003 1.000 1.015 1 0.981
-security_pgp_d 1.861 0.993 1.000 0.991 13.144 1
-security_pgp_e 1.861 1.017 1.000 1.006 13.144 1.006
-security_rijndael_d 1.000 0.990 3.823 0.966 1 0.966
-security_rijndael_e 1.000 1.001 3.823 0.974 1 0.995
-security_sha 0.999 1.270 1.000 1.174 0.999 1.501
-telecom_adpcm_c 1.000 1.000 1.000 1.008 1 1.038
-telecom_adpcm_d 1.000 0.997 1.000 1.132 1 1.138
-telecom_CRC32 1.000 1.094 1.000 1.042 1 1.112
-telecom_gsm 1.810 1.076 1.815 1.15 1.81 1.117
-Geomean 1.152 1.021 1.818 1.01 2.442 1.024
-The results are shown in Tables 7 and 8.
-We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively.
-LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Module and Function level for Inline and Loop level for Loop Unroll, is outside the scope of this work.
-Instead, we showcase the overall benefits of each benchmark.
-For instance, Table 8 showcases the performance gains using ACPO-FI, ACPO-LU, and the combined version with separate columns.
+Table 2. ACPO Experimental Results (Polybench)
+Benchmark ACPO-LU ACPO Combined
+Size Bloat % Speedup Size Bloat % Speedup
+2mm 1.903 0.984 1.903 0.984
+3mm 1.904 0.983 2.805 0.976
+adi 1.0002 1.012 1.38 1.04
+atax 0.999 0.985 1 0.987
+bicg 1 1.02 1 1.022
+cholesky 1.001 1.102 1 1.063
+correlation 1.005 0.957 1.901 1.002
+covariance 1.001 1.003 1.902 1.009
+deriche 1.0018 1.078 1.902 1.115
+doitgen 1.0001 1.011 1.901 1.013
+floyd-Warshall 1.0004 1.493 1 0.974
+gemver 1.0004 0.958 1.902 1.494
+gesummv 1 1.143 1 0.975
+gramschmidt 1 0.91 1 0.995
+lu 1.001 0.978 1.9 1
+ludcmp 1.0008 1.047 1 0.996
+mvt 1 1.046 1 1.066
+nussinov 2.803 1.171 1 1.013
+seidel-2d 1.0005 1.0008 1 1.174
+symm 1.0004 1.06 1.902 1.072
+syr2k 1.0009 0.991 1.903 0.989
+syrk 1.0005 1.002 1.903 1.036
+trisolv 1 1.223 1 1.232
+trmm 1.0004 0.946 1 0.975
+Geomean 1.10193 1.040725 1.383 1.045
 1//Security_sha Compilation (ACPO)
@@
 Listing 6: O3 Decisions
+Table 3. ACPO Experimental Result (Cbench)
+Sz%: Size bloat, Sp: Speedup. Multicolumns 1, 3, and 4 are wrt. O3 and 2, is ACPO-FI compared against MLGO (Trofin et al., 2021).
+Benchmark ACPO-FI wrt. MLGO ACPO-LU Combined
+SzSp SzSp SzSp SzSp
+auto_bitcount 1 0.999 1 1.005 1.001 1.003 1 1.001
+auto_qsort1 1 0.994 1 1.003 1 0.982 1.91 0.989
+auto_susan_c 1 0.999 1.567 1.122 3.685 1.035 19.797 1.066
+auto_susan_e 1 1.023 1.567 1.04 3.685 1.019 19.797 1.026
+auto_susan_s 1 0.98 1.567 1.008 3.685 1.038 19.797 0.995
+bzip2d 1.441 0.991 1.239 1.021 2.224 1.01 3.541 1.014
+bzip2e 1.441 1.007 1.189 1.041 2.35 1.001 3.546 1.015
+cons_jpeg_c 1.288 1.017 1.2 1.005 5.385 0.986 5.086 1.027
+cons_jpeg_d 1.289 0.923 1.19 1.335 5.4 1.073 4.807 1.026
+cons_lame 1.437 1.027 1.137 1.007 4.501 1.03 4.729 1.034
+cons_mad 1.217 1.071 1.16 1 4.321 0.99 3.193 0.968
+cons_tiff2bw 1.18 1.002 1.351 1.23 4.502 0.943 2.471 0.987
+cons_tiff2rgba 1.18 1.25 1.356 1.11 4.681 1.036 2.657 1.012
+cons_tiffdither 1.18 1.046 1.14 1.014 1.213 0.967 2.655 1.002
+cons_tiffmedian 1.18 1.017 1.048 1.056 1 0.97 3.021 0.983
+network_dijkstra 1 0.993 1 0.996 1 0.99 1 1.037
+network_patricia 1 0.98 1.052 0.99 1 0.974 1 1.004
+office_ghostscript 1.275 1 1.16 1.04 1 1 1.12 1.001
+office_ispell 0.996 1 1.17 1.02 1 1 1 1
+office_rsynth 1 1 1.12 1.04 2.024 1.01 2.024 1.011
+office_strings 1 1.006 1 1 1 1 1 1
+secur_blowfish_d 1 1.002 1 1.002 1 0.978 1 0.965
+secur_blowfish_e 1 1.003 1.39 1.029 1 1.015 1 0.981
+security_pgp_d 1.861 0.993 1.12 1 1 0.991 13.144 1
+security_pgp_e 1.861 1.017 1.12 1.042 1 1.006 13.144 1.006
+secur_rijndael_d 1 0.99 1.39 1.039 3.823 0.966 1 0.966
+secur_rijndael_e 1 1.001 1.09 1.035 3.823 0.974 1 0.995
+security_sha 0.999 1.27 1.009 1.132 1 1.174 0.999 1.501
+tele_adpcm_c 1 1 0.996 1.001 1 1.008 1 1.038
+tele_adpcm_d 1 0.997 1 1 1 1.132 1 1.138
+telecom_CRC32 1 1.094 1 1.066 1 1.042 1 1.112
+telecom_gsm 1.81 1.076 1.05 1.04 1.815 1.15 1.81 1.117
+Geomean 1.152 1.021 1.161 1.041 1.818 1.01 2.442 1.024
+The combined results are shown in Tables 2 and 3.
+We observe a slight increase in performance benefit for both benchmarks, i.e., from 4% to 4.5% for Polybench and from 2% to 2.4% for Cbench, respectively.
+Note that we don’t deploy ACPO-FI on Polybench as it consists of small nested-loop kernels which might not directly benefit from an optimized inliner, however, the experimental result reveals that indirectly it can.
+The underlying reason is that by it can provide more opportunities for the subsequent passes registered under O3 (Ashouri et al., 2022).
+LLVM’s O3 pipeline has a complex behavior when it comes to pass activation and as mentioned earlier, several passes including Loop Unrolling are activated multiple times thus, it is a hard task to quantify the benefits of each model when deployed together and the accurate measurement of partial benefits, especially when optimizations are working at different scopes, i.e., Call Graph vs Function level.
 One exceptionally good performing benchmark is security_sha, which we had already optimized by 27% (show on Table 3) using ACPO’s Inliner, now with 50% performance gain.
@@
 Further inspection revealed that O3’s threshold mechanism built into the current version of Loop Unroll pass was limiting the assignment of higher values of unroll count even though the trip count was available at compile time.
-Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism as it is less than or equal to the calculated cost.
-This example showcases the superiority of using 30/13 well-representative loop/function features we designed in training ACPOLUModels to deduce the kind of complex correlation between a code segment and the value of unroll count and the Boolean decision of inlining a callsite to optimize the performance.
+Similarly, the decision to not inline several callsites including sha_transform is limited by the Inliner’s threshold mechanism.
 6. Discussion
-ACPO is the first scalable framework for bringing ML into compilers.
-It provides the ability to reproduce models on different architecture very fast and the methodology lays out the steps to regenerate data, retrain a new model, and easily deploy with LLVM.
-Compiler autotuning and optimization space have benefited from ML for almost two decades.
-As we pointed out earlier, there are many recent works leveraging ML, however, the goal of ACPO is to provide a novel flexible framework for the LLVM community to leverage ML on any optimization passes of interest with minimal efforts.
-Nevertheless, there are a number of challenges and shortcomings we would like to mention here.
-Integration and dependencies
-ACPO attempts to alleviate the challenging task of supply chain management and the use of licensed tools by separating the ML framework from the compiler.
-We showed both the integration of PyTorch and TensorFlow among other moving objects, i.e., Autotuner, into our framework, however, upstreaming would remain challenging and requires community efforts to shape the next generations of ML-enabled compilers.
-We also provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers.
+ACPO provides a class of features at varying LLVM scopes, therefore, new passes can utilize the already available features with minimal or no efforts.
+It also provides the flexibility of supporting other types of features, i.e., Graph-based, embeddings, etc.
+Additionally, we provide the Ahead-Of-Time (AOT) compilation facilities as an alternative method of inference to the developers.
 As mentioned in Section 3.5, ACPO uses Python named pipes to perform interprocess communication.
@@
 The flexibility of choosing between the two modes allows for a fast iterative prototyping of ML models during design and training without the need to rebuild the compiler infrastructure after every change and also, an efficient compilation overhead at deployment once the ML models are optimized and ready to be shipped.
-Multi-objective optimization
-ACPO optimizes the performance of a code region with intelligent unrolling decisions.
-However, in a number of cases, we experimentally observe that the emitted code size is also increased.
-This is a(n) (in)direct result of Inliner and Loop Unroll which increase the number of instructions and thus, the code size.
-Although this is outside the scope of the current work, an enhanced version may also benefit from taking into account both of the objectives in its exploration strategy by identifying the Pareto optimality (Hoste and Eeckhout, 2008; Ashouri et al., 2013; Moreau et al., 2018; Silva et al., 2021).
 7. Conclusion
 In this paper, we presented ACPO, the first scalable and modular ML-guided compiler framework capable of optimizing performance for LLVM.
-We demonstrated that using ACPO framework, we are able to enable Loop Unroll pass to benefit from a seamless integration of an ML model, i.e., ACPOLUModel, for which we experimentally achieve on average a 4%, 3%, 5.4%, and 0.2% performance speedup over LLVM’s O3 on Polybench, Coral-2, CoreMark, and Graph-500, respectively.
-Additionally, we showcased a scenario where we leverage an existing ML model to optimize Function Inlining pass and similar to the original model’s performance, we experimentally obtained 2% speedup wrt LLVM’s O3 when deployed on CBench and we showed that by combining both models, the benefits increase up to 4.5% and 2.4% on Polybench and Cbench, respectively.
-Future works will be focused on extending ACPO framework by enabling a number of other optimization passes inside LLVM’s pipeline and leveraging innovative software characterization methods to be employed.
-Appendix A Appendix —  ACPO Headers
-1class ACPOModel {
-2public:
-3  ACPOModel(...);
-4  ~ACPOModel();
-5  void setMLIF(std::shared_ptr<ACPOMLInterface> ML);
-6  std::shared_ptr<ACPOMLInterface> getMLIF();
-7  void addRequiredResultField(std::string Name, Type::TypeID &ID);
-8  std::unique_ptr<ACPOAdvice> getAdvice();
-9  void addFeature(int64_t ID, Constant *Val);
-10  void setCustomFeatures(const std::vector<std::pair<std::string, Constant *>> &FeatureValues);
-11
-12protected:
-13  virtual void prepareModelInput();
-14  virtual bool runModel(std::unique_ptr<ACPOAdvice> &Result);
-15};
-Listing 7: ACPOModel Header
-1class ACPOMLInterface {
-2public:
-3  ACPOMLInterface();
-4  virtual ~ACPOMLInterface();
-5  virtual bool loadModel(std::string ModelSpecFile);
-6  virtual bool freeModel(std::string ModelName);
-7  virtual bool registerModel(std::string ModelName, int NumFeatures, int NumOutputs);
-8  virtual bool registerFeature(std::string ModelName, std::string FeatureName, int Index);
-9  virtual bool registerOutput(std::string ModelName, std::string OutputName, std::string OutputType);
-10  virtual bool setCustomFeature(std::string ModelName, uint64_t FeatureID, int FeatureValue);
-11  ...
-12  virtual bool setCustomFeature(std::string ModelName, std::string FeatureName, bool FeatureValue);
-13  virtual bool setCustomFeatures(std::string ModelName, const std::vector<std::pair<uint64_t, std::string>> &FeatureValues);
-14  virtual bool setCustomFeatures(std::string ModelName, const std::vector<std::pair<std::string, std::string>> &FeatureValues);
-15  ...
-16  virtual bool runModel(std::string ModelName);
-17  virtual std::string getOutputType(std::string ModelName, std::string OutputName);
-18  virtual int getModelResultI(std::string OutputName);
-19  ...
-20  virtual bool getModelResultB(std::string OutputName);
-21  ...
-22  virtual int getStatus();
-23  virtual bool closeMLInterface();
-24};
-Listing 8: ACPOMLInterface Header
+We demonstrated that using ACPO framework, we are able to enable Loop Unroll and Function Inline passes to benefit from a seamless integration of ML models, i.e., ACPOLUModel and ACPOFIModel, for which we experimentally achieved on average 4% and %4.5 on Polybench, and 2.1% and 2.5% on Cbench.
+ACPO provides a generalized feature collection class for which compiler developers can add to their instantiated ACPOModel class of their need and provides utility classes to use other types of features as well.
+Future works will be focused on extending ACPO framework by enabling other features and optimization passes inside LLVM’s pipeline and providing innovative software characterization methods using large Languages Models (LLMs).